# Mixed-Integer Programming STALC Stratified Topological Autonomy for Long-Range Coordination

### v1

@@ IX-D3 Urban Environment Experiments
We generated graph plans for the segmentations of the environment shown in Fig. 20(a) and Fig. 20(b) using two different scales for overwatch, as described in Sec. V-C. A particular scenario may consider overwatch to be more valuable and the scale factor allows that high-level guidance to vary the dynamic topological graph we consider. This demonstrates the control the user has over the high-level behaviors of the robots when generating these highly complicated paths. The resulting graph plans are shown in Fig. 21 and Fig. 22, respectively. In Fig. 23 we do not scale the overwatch (i.e., scale of 1) to instead demonstrate planning with two observer locations. From the observer regions indicated in Fig. 20, Fig. 24 shows segments of the observer’s view when the robots were able to be detected in Urban Graphs 1 and 2.
Fig. 1 shows a third person view of the robots arranged at node 4 before moving into formation to traverse edge (4,3) in the first step of the plan in Fig. 21(a). Fig. 1 denotes the leader and follower goal points in cover region 5 and shows the MPPI trajectories of the two robots that begin moving first. The third robot waits for the other two to get ahead of it before moving into it’s position in the formation. The included video shows the STALC algorithm running in this urban environment example following the paths in Fig. 21(a).
(a) Urban Graph 1 robots moving in formation on edge (4,3).
(b) Urban Graph 1 robot on edge (2,0).
(c) Urban Graph 2 robot on edge (5,3).
Figure 24: Observer views for Urban Graphs 1 and 2. In (c), the tower the observer is on in Urban Graph 1 is shown on the right.
@@ X Conclusion
Our hierarchical planning approach is able to effectively generate coordinated tactics using a priori data while compensating for changing environments with the lower-levels of the autonomy stack. We demonstrated expressing complicated scenarios compactly with a dynamic topological graph and efficient MIP formulation to yield rapid solutions to real world scenarios. We evaluated our approach in simulation and hardware experiments showing the full system operation, including swapping between overwatch and formation conditions in highly constrained environments with static and dynamic obstacles. Future research will include receding-horizon graph-based planning to update the visibility map online to account for changing conditions and a mid-range planner that utilizes the real-time sensor readings of the environment.
@@ Acknowledgment
We gratefully acknowledge the support of the Army Research Laboratory under grant W911NF-22-2-0241.

### v2

@@ IX-D3 Urban Environment Experiments
We generated graph plans for the segmentations of the environment shown in Fig. 21(a) and Fig. 21(b) using two different scale factors for overwatch. As described in Sec. VI-C, a particular scenario may consider overwatch to be more valuable and the overwatch scale factor allows that high-level guidance to vary the DTG. This graph variation demonstrates the control the user has over the high-level behaviors of the robots when generating these coordinated maneuvers. The resulting graph plans are shown in Fig. 22 and Fig. 23. In Fig. 24 we do not scale the overwatch (i.e., scale of 1) to instead demonstrate planning with two observer locations. Fig. 25 shows segments of the observers’ views when the robots were able to be detected in Urban Graphs 1 and 2.
Fig. 1 shows a third person view of the robots arranged at node 4 before moving into formation to traverse edge (4,3) in the first step of the plan in Fig. 22(a). Fig. 1 denotes the leader and follower goal points in cover region 5 and shows the MPPI trajectories of the two robots that begin moving first. The third robot waits for the other two to get ahead of it before moving into its position in the formation.
(a) Urban Graph 1: robots moving in formation on edge (4,3).
(b) Urban Graph 1: robot on edge (2,0).
(c) Urban Graph 2: robot on edge (5,3).
Figure 25: Observer views for Urban Graphs 1 and 2. The traversing robots are circled in red. In (c), the tower the observer is on in Urban Graph 1 is shown on the right.
@@ X Conclusion
Our hierarchical planning approach STALC is able to efficiently generate coordinated tactics using a priori data while compensating for changing environments with the lower-levels of the autonomy stack. We demonstrated expressing complicated scenarios compactly with a Dynamic Topological Graph (DTG) and efficient MILP formulation to yield rapid solutions for end-to-end plans in real-world scenarios. We evaluated our approach in simulation and hardware experiments showing the full system’s autonomous operation, including swapping between overwatch and formation conditions in highly constrained environments with static and dynamic obstacles. Future research will include receding-horizon, graph-based planning to update the visibility map online to account for changing conditions, and a mid-range planner that utilizes real-time sensor readings of the environment.
@@ Acknowledgment
Research was sponsored by the Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-22-2-0241 and W911NF-25-2-0033. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.